# PatchAD: Patch-based MLP-Mixer for Time Series Anomaly Detection

### v3

@@ 4.3.4 Time Cost
We conducted a time consumption analysis on the SMD dataset under identical experimental conditions. Figure 3(b) demonstrates that under different sliding step conditions, our method outperforms DCdetector. Specifically, when the step is set to 1, PatchAD’s speed is 1.45x faster than DCdetector. Additionally, Figure 3(d) illustrates the time required for model training with a fixed window size (window size=105). This indicates that PatchAD exhibits higher efficiency in training compared to DCdetector. Figure 3(c) shows the time required for model training with a fixed sliding step (step=35) while varying the window size from 35 to 175 for different $d_{model}$ sizes. Both graphs show that the time consumption of PatchAD exhibits a linear growth with respect to the independent variables. Thus, increasing the window size or enlarging $d_{model}$ has a relatively minor impact on PatchAD. However, the time step noticeably affects the duration as this parameter leads to increased data volume.
@@ 5 Conclusion
We presents PatchAD, an innovative algorithm tailored for time series anomaly detection. PatchAD developed a multi-scale, patch-based MLP Mixer architecture grounded in contrastive learning. This architecture is unique in its use of a purely MLP-based approach to discern both inter-patch and intra-patch relationships in time series sequences. By amplifying the differences between two data views, PatchAD enhance its capability to differentiate between normal and anomalous patterns. Furthermore, PatchAD adeptly learns the relationships between different data channels and incorporates a Dual Project Constraint to avert model degradation. Moreover, the efficacy of PatchAD, including its various components, is validated through extensive experimentation. The results demonstrate that proposed PatchAD outperforms other methods across eight benchmark datasets.
@@
This is the appendix of PatchAD: Patch-based MLP-Mixer for Time Series Anomaly Detection.
Datasets
Many publicly available real-world datasets are extensively used in existing methods. We utilize eight datasets to evaluate model performance: (1) MSL (Mars Science Laboratory dataset), (2) SMAP (Soil Moisture Active Passive dataset), (3) PSM (Pooled Server Metrics), (4) SMD (Server Machine Dataset), (5) SWaT (Secure Water Treatment Testbed), (6) WADI (Water Distribution Testbed), (7) NIPS-TS-SWAN, and (8) NIPSTS-GECCO. Table 4 shows the details.
@@
Baslines
We compared our model against 27 benchmark models, including: (1) Reconstruction-based: AutoEncoder Sakurada and Yairi [2014], LSTM-VAE Park et al. [2018], OmniAnomaly Su et al. [2019], BeatGAN Zhou et al. [2019], InterFusion Li et al. [2021b]; (2) Autoregression-based: VAR Anderson [1976], Autoregression Rousseeuw and Leroy [2005], LSTM-RNN Bontemps et al. [2016], LSTM Hundman et al. [2018], CL-MPPCA Tariq et al. [2019]; (3) Density-based: LOF Breunig et al. [2000], MMPCACD Yairi et al. [2017], DAGMM Zong et al. [2018]; (4) Clustering-based: DeepSVDD Ruff et al. [2018], THOC Shen et al. [2020], ITAD Shin et al. [2020]; Matrix Profile Yeh et al. [2016]; (5) The classic methods: OCSVM Schölkopf et al. [2001], OCSVM-based subsequence clustering (OCSVM*), IForest Liu et al. [2008], IForest-based subsequence clustering (IForest*), Gradient boosting regression (GBRT) Elsayed et al. [2021]; (6) Change point detection and time series segmentation methods: BOCPD Adams and MacKay [2007], U-Time Perslev et al. [2019], TS-CP2 Deldari et al. [2021]; (7) Transformer-based: Anomaly Transformer  [^7] https://github.com/thuml/Anomaly-Transformer Xu et al. [2021]; DCdetector  [^8] https://github.com/DAMO-DI-ML/KDD2023-DCdetector Yang et al. [2023]; These two are current state-of-the-art methods.
Implementation Details
All experiments were conducted using PyTorch implementation on a single NVIDIA GTX 1080TI GPU. We summarize the default hyperparameters of PatchAD as follows: PatchAD consists of 3 encoder layers ($L$=3). The hidden layer dimension $d_{{model}}$ is set to 40. The default projector constraint coefficient is $0.2$. The Adam optimizer’s default learning rate is $10^{-4}$, with a batch size of 128, trained for 3 epochs across all datasets. Different patch sizes and window sizes were chosen for each dataset, defaulting to [3, 5] and 105, respectively. If an anomaly score at any given timestamp exceeds a threshold $\sigma$, the model identifies it as an anomaly. The default value for $\sigma$ is 1. For details on parameter selection, please refer to the Appendix.
Algorithm’s Pseudo-code
@@
final_loss = (1-c) * loss1 + c * loss_proj
@@
Full Comparison Results
In Table 5, Aff-P and Aff-R are the precision and recall of affiliation metric Huet et al. [2022], respectively. R_A_R and R_A_P are Range-AUC-ROC and Range-AUC-PR Paparrizos et al. [2022], respectively. V_ROC and V_RR are volumes under the surfaces created based on ROC curve and PR curve Paparrizos et al. [2022], respectively. Besides, we give more results in table 6, and the results from more real-world multivariate time-series datasets show that our proposed PatchAD has good robustness as well as performance.
@@
Table 5: Multi-metrics results on all real-world multivariate datasets. All results are in %, the best ones are in Bold, and the second ones are underlined.
@@
Table 6: More results on all real-world multivariate datasets. All results are in %, the best ones are in Bold, and the second ones are underlined.

### v5

@@ 4.4.4 Time Cost
We conducted a time consumption analysis on the SMD dataset under identical experimental conditions. Figure 5(b) demonstrates that under different sliding step conditions, our method outperforms DCdetector. Specifically, when the step is set to 1, PatchAD’s speed is 1.45x faster than DCdetector. Additionally, Figure 5(d) illustrates the time required for model training with a fixed window size (window size=105). This indicates that PatchAD exhibits higher efficiency in training compared to DCdetector. Figure 5(c) shows the time required for model training with a fixed sliding step (step=35) while varying the window size from 35 to 175 for different $d_{model}$ sizes. Both graphs show that the time consumption of PatchAD exhibits a linear growth with respect to the independent variables. Thus, increasing the window size or enlarging $d_{model}$ has a relatively minor impact on PatchAD. However, the time step noticeably affects the duration as this parameter increases data volume.
4.4.5 Model Size
Figure 2 compares the sizes of different models on SWAT dataset. Our model has a compact size of only 3.2MB, which is smaller than both AnomTrans and DCdetector, and significantly smaller than the LLM model (GPT2-Adapter) and diffusion model (D3R). This observation indicates that our model is more lightweight and resource-efficient.
@@ 5 Conclusion
We presents PatchAD, an innovative algorithm tailored for time series anomaly detection. PatchAD developed a lightweight multi-scale, patch-based MLP Mixer architecture grounded in contrastive learning. This architecture is unique in using a purely MLP-based approach to discern inter-patch and intra-patch relationships in time series sequences. By amplifying the differences between two data views, PatchAD enhance its capability to differentiate between normal and anomalous patterns. Furthermore, PatchAD adeptly learns the relationships between different data channels and incorporates a Dual Project Constraint to avert model degradation. Moreover, the efficacy of PatchAD, including its various components, is validated through extensive experimentation. The results demonstrate that proposed lightweight PatchAD outperforms over 30 methods across nine benchmark datasets from different application scenarios.
@@
This is the appendix of PatchAD: A Lightweight Patch-Based MLP-Mixer for Time Series Anomaly Detection.
Appendix A Theory
In this section, our objective is to perform a theoretical analysis of PatchAD’s effectiveness. We want to prove that normal time series data points exhibit reduced information entropy. We have made the following assumptions based on the theoretical analysis presented in previous studies.
Assumption 1: The distribution of normal data differs from that of anomalies, and both consist of two components, denoted as $\tau+\alpha$. Here, $\tau$ represents the underlying data pattern, while $\alpha$ represents the pattern of normal or anomalous data. Therefore, we denote normal data distribution by $\Pi$ and anomalies as $\Omega$. As a result, the normal data $\mathcal{X}_{norm}$ and the anomalous data $\mathcal{X}_{anom}$ can be expressed as follows:
$\mathcal{X}_{norm}=\tau+\pi,\text{ and }\mathcal{X}_{anom}=\tau+\omega.$ (12)
Assumption 2: The number of normal data samples in the dataset is significantly larger than that of anomalous data samples. Therefore, we have the following relationship:
$|\mathcal{X}_{norm}|=\zeta\cdot|\mathcal{X}_{anom}|>|\mathcal{X}_{anom}|,$ (13)
where ||$\cdot$|| ⋅ | represents the number of samples in the data, and $\zeta$ is a large number. We denote the entire dataset as $\mathcal{D}=\{\mathcal{X}_{norm},\mathcal{X}_{anom}\}$.
Proof 1: The entropy of the anomalous data is higher than that of normal data. For a random variable $X$, its entropy is defined as, and the entropy serves as a metric for quantifying the uncertainty and complexity of the data [54]:
$\operatorname{H}(X)=-\sum_{i}p(x_{i})\log_{2}p(x_{i})$ (14)
[…]
Using the addition rule of probability, we have $p(\tau+\pi)=p(\tau)\cdot p(\pi)$. Substituting this into the previous equation, we obtain: $\operatorname{H}(\mathcal{X}_{norm})=-\int p(\tau)\cdot p(\pi)\log(p(\tau)\cdot p(\pi))d(\tau+\pi).$ Expanding further, we can deduce:
$\operatorname{H}(\mathcal{X}_{norm})=-\int p(\tau)\cdot p(\pi)(\log p(\tau)+\log p(\pi))d(\tau+\pi).$ (15)
[…]
Due to the significantly larger number of samples in normal data compared to anomalous data, i.e., $|\mathcal{X}_{norm}|=\zeta\cdot|\mathcal{X}_{anom}|>|\mathcal{X}_{anom}|$, we can extract the large number $\zeta$ and use the properties of logarithm to move it inside the $\log$ function:
$\operatorname{H}(\mathcal{X}_{norm})=-\zeta\sum_{x\in\mathcal{X}_{norm}}P^{\prime}(x)\log P^{\prime}(x)-\zeta\sum_{x\in\mathcal{X}_{norm}}P^{\prime}(x)\log\zeta$ (16)
The second term $-\zeta\sum_{x\in\mathcal{X}_{norm}}P^{\prime}(x)\log\zeta$ is a constant term, denoted as $C$.
Therefore, we can express the entropy of normal data as:
$\operatorname{H}(\mathcal{X}_{norm})=-\zeta\sum_{x\in\mathcal{X}_{norm}}P^{\prime}(x)\log P^{\prime}(x)+C.$ (17)
Similarly, we can derive the entropy of anomalous data $\operatorname{H}(\mathcal{X}_{anom})$:
$\operatorname{H}(\mathcal{X}_{anom})=-\sum_{x\in\mathcal{X}_{anom}}P^{\prime}(x)\log P^{\prime}(x).$ (18)
At this point, we can conclude that $\operatorname{H}(\pi)<\operatorname{H}(\omega)$. Therefore, the entropy of normal data $\operatorname{H}(\mathcal{X}_{norm})$ is smaller than the entropy of anomalous data $\operatorname{H}(\mathcal{X}_{anom})$.
Appendix B Experimental Settings
B.1 Datasets
We applied eight multivariable and one univariable datasets for evaluation: (1) MSL (2) SMAP  [^2] MSL & SMAP: https://github.com/ML4ITS/mtad-gat-pytorch(3) PSM (4) SMD  [^3] https://github.com/NetManAIOps/OmniAnomaly (5) SWaT  [^4] https://itrust.sutd.edu.sg/itrust-labs_datasets (6) WADI  [^5] https://itrust.sutd.edu.sg/testbeds/water-distribution-wadi (7) NIPS-TS-SWAN (8) NIPSTS-GECCO  [^6] Others:http://github.com/DAMO-DI-ML/KDD2023-DCdetector. (9) UCR  [^7] https://www.cs.ucr.edu/~eamonn/time_series_data_2018/UCR_TimeSeriesAnomalyDatasets2021.zip Table 4 shows the details.
@@ B.1 Datasets
B.2 Baslines
We compared our model against 35 benchmark models, including: (1) Reconstruction-based: AutoEncoder [55], LSTM-VAE [6], OmniAnomaly [7], BeatGAN [43], InterFusion [47], USAD [17], COUTA[24]; (2) Autoregression-based: VAR [12], Autoregression [56], LSTM-RNN [13], LSTM [46], CL-MPPCA [41], TimesNet [14], NPSR-pt and NPSR-seq [50]; (3) Density-based: LOF [8], MMPCACD [40], DAGMM [9]; (4) Clustering-based: DeepSVDD [45], THOC [48], ITAD [39]; Matrix Profile [57]; (5) The classic methods: OCSVM [19], OCSVM-based subsequence clustering (OCSVM*), IForest [37], IForest-based subsequence clustering (IForest*), Gradient boosting regression (GBRT) [58]; (6) Change point detection and time series segmentation methods: BOCPD [44], U-Time [38], TS-CP2 [42]; (7) Transformer-based: Anomaly Transformer  [^8] https://github.com/thuml/Anomaly-Transformer [27], DCdetector  [^9] https://github.com/DAMO-DI-ML/KDD2023-DCdetector [10], TranAD [59]; (8) Large Language Model: GPT2-Adapter[15]; (9) Diffusion model: D3R [60]; These two are current state-of-the-art methods.
B.3 Implementation Details
All experiments used PyTorch implementation on a single NVIDIA GTX 1080TI GPU. We summarize the default hyperparameters of PatchAD as follows: PatchAD consists of 3 encoder layers ($L$=3). The hidden layer dimension $d_{{model}}$ is set to 40. The default projector constraint coefficient is $0.2$. The Adam optimizer’s default learning rate is $10^{-4}$, with a batch size of 128, trained for 3 epochs across all datasets. Different patch sizes and window sizes were chosen for each dataset, defaulting to [3, 5] and 105, respectively. If an anomaly score of any given timestamp exceeds a threshold $\sigma$, the model identifies it as an anomaly. The default value for $\sigma$ is 1.
Appendix C Algorithm’s Pseudo-code
@@
loss_rec = MSE(rec_x, x)
final_loss = (1-c) * loss1 + c * loss_proj + loss_rec
@@
Appendix D Unveiling the Mechanisms
Figure 8 illustrates the knowledge acquired by PatchAD in four distinct cases. Cases 2, 3, 4 and 5 have a patch size of 1, whereas case 1 has a patch size of $[1,2,5]$. The "Original" represents the original data, while "Rec 1" and "Rec 2" correspond to the features learned by the Inter Mixer and the Intra Mixer, respectively. It is observed that the Intra Mixer captures the generalized knowledge in the data, which encompasses the overall trend information. Conversely, the Inter Mixer learnsthe detailed features of the data, compensating for any aspects that may not be captured by the Intra Mixer. A comparison between case 1 and case 2 reveals that using three distinct patch scales facilitates the acquisition of more detailed representations and enhances the generalization of the acquired features. Note that the y-axis ranges for Figure 8(a) and Figure 8(b) are different. Moreover, these two mixers collaborate to segregate and amplify the distinctive features effectively. The decoupling of features assists PatchAD in amplifying the anomaly-related characteristics, as indicated in Figure 3(c), thereby enhancing the detectability of anomalies. The superior performance of PatchAD can also be attributed to the varying sizes of the Inter and Intra Mixer branches, which prevent overfitting. This is the reason why PatchAD does not require the construction of negative samples in contrastive learning, as the two different networks construct implicit positive and negative samples.
Moreover, we empirically validate this conclusion through experiments. Case 1 was chosen as the baseline, and we investigated the entropy of the features acquired by the Inter and Intra Mixers under varying patch size conditions. Figures 9(a) and 9(b) illustrate that both the Inter Mixer and Intra Mixer display a consistent trend of decreasing entropy throughout the learning process. This observation suggests that the models effectively acquire knowledge and reduce uncertainty. Furthermore, we noticed that the Inter Mixer demonstrates the highest learning speed and exhibits the most substantial decrease in entropy when the patch size is configured as [1, 2, 5]. Additionally, as the patch size decreases, the entropy of the Inter Mixer increases, and it actually rises when the patch size is set to 1. This implies that the network is required to learn an excessive number of features. Figure 9(c) computes the ratio of Intra to Inter entropy, and our findings indicate that as the patch size increases, the features acquired by the Intra Mixer gain greater significance, thereby accelerating model learning through the utilization of multi-scale patch sizes.
Figure 10 illustrates the visualization of feature distributions in the SWAT dataset under three conditions. It is observed that anomalies are more easily detected in the Intra Mixer view (See Figure 10(b) ), as its distinctiveness is more pronounced compared to the other two views (Figure 10 (a) & (c) ).
According to our analysis, maximizing the disparities between the Inter Mixer and Intra Mixer can enhance the model’s detection capabilities. During network optimization, the Inter-Intra loss compels the Intra Mixer to acquire more generalized representations, effectively mitigating model overfitting.
Figure 8: Visualization of the features learnt by Inter and Intra Mixer.
Figure 9: Entropy analysis.
Figure 10: Latent distribution’s visualization by T-SNE.
Appendix E Parameter Sensitivity
We conducted a sensitivity analysis on PatchAD’s parameters. The default parameter settings remain consistent with the basic configuration. Figure 6(a) demonstrates the impact of the projection head constraint on PatchAD. Overall, it holds more significance for smaller-scale datasets like MSL. However, excessively strong constraints hinder the model’s learning process. Therefore, setting this constraint between 0.1 to 0.4 is more advisable. Figure 6(b) analyzes PatchAD’s final model performance when the anomaly threshold $\sigma$ ranges from 0.5 to 1.2. The selection of the anomaly threshold appears more robust on the PSM and SMAP datasets than the MSL dataset. Setting this value between 0.8 and 1 for most datasets yields optimal results. Figure 6(c) illustrates how the model’s performance is influenced by the network dimension $d_{model}$. PatchAD demonstrates better results with smaller $d_{model}$ sizes. Therefore, to balance model performance and complexity, we set $d_{model}=40$ to ensure PatchAD operates effectively. Figure 6(d) indicates that PatchAD’s performance is influenced by the number of encoder layers, and given a compromise, we set it to 3 or 4. Figure 6(e) demonstrates PatchAD’s robustness concerning the window size, which ranges from 60 to 210. The time window is a critical parameter, typically set to an initial value of 105 across most datasets. Lastly, Figure 6(f) displays PatchAD’s performance under different multi-scale patch size combinations. (e.g., a patch size of [3,5] represents using two scales in PatchAD.) The outcomes suggest that combining various patch sizes can enhance PatchAD’s performance.
Appendix F Visualization comparison
Figure 11 shows the performance of WindowAD algorithms in the face of different types of anomalies. These methods have difficulty in detecting contextual and group point anomalies and are able to detect seasonal anomalies, but there are some false positives, which indicates that it has a low precision.
(a) Contextual
(b) Seasonal
(c) Group
Figure 11: Visualization of WindowAD-based algorithm’s score.
Appendix G Full Comparison Results
Table 6 presents the performance of different methods on four datasets, namely WADI, SWAT, PSM, and SWAN. This is the original data of Figure2. We have selected representative algorithms for experimental comparison.
@@ Appendix G Full Comparison Results
Table 5: Overall results on NIPS-TS datasets. All results are in %, the best in Bold, and the second in underlined.
Dataset NIPS-TS-SWAN NIPS-TS-GECCO Average F1
Metric P R F1 P R F1
MatrixProfile [57] 16.70 17.50 17.10 4.60 18.50 7.40 12.25
GBRT [58] 44.70 37.50 40.80 17.50 14.00 15.60 28.20
LSTM-RNN [13] 52.70 22.10 31.20 34.30 27.50 30.50 30.85
Autoregression [56] 42.10 35.40 38.50 39.20 31.40 34.90 36.70
OCSVM [19] 47.40 49.80 48.50 18.50 74.30 29.60 39.05
IForest* [37] 40.60 42.50 41.60 39.20 31.50 39.00 40.30
AutoEncoder [55] 49.70 52.20 50.90 42.40 34.00 37.70 44.30
AnomalyTrans [27] 90.70 47.40 62.30 25.70 28.50 27.00 44.65
IForest [37] 56.90 59.80 58.30 43.90 35.30 39.10 48.70
AnomalyTrans [27] 90.71 47.43 62.29 29.96 48.63 37.08 49.69
DCdetector [10] 96.57 59.08 73.31 38.41 59.73 46.76 60.04
PatchAD (Ours) 96.75 59.15 73.41 40.62 62.88 49.35 61.38
Table 6: The performance of algorithms on different evaluation metrics without point adjustment.
Dataset WADI SWAT
Metric P R F1_ori Aff-Pre AFF-Rec Aff-F1 AUC P R F1_ori Aff-Pre AFF-Rec Aff-F1 AUC
Deep SVDD 4.05 38.53 7.34 50.06 55.79 52.77 40.53 12.74 98.50 22.57 53.58 99.99 69.77 86.81
USAD 5.74 100.00 10.86 54.92 100.00 70.90 46.39 12.17 100.00 21.70 53.00 100.00 69.28 88.66
COUTA 98.72 15.58 26.92 99.18 23.08 37.44 49.48 94.25 6.33 11.87 79.98 33.20 46.92 75.54
TranAD 6.49 63.89 11.78 55.91 91.96 69.54 52.60 14.70 96.40 25.50 55.65 99.66 71.42 88.90
TimesNet 64.04 10.24 17.65 65.25 39.30 49.05 65.08 11.05 3.21 4.98 57.84 93.77 71.55 26.76
AnomTrans 4.14 14.12 6.40 45.73 40.66 43.05 49.20 4.81 29.32 8.27 53.03 98.08 36.08 19.20
DCdetector 5.80 14.42 8.27 59.63 90.98 69.98 49.88 15.49 5.54 8.16 52.76 97.33 68.94 50.68
GPT2-Adapter 5.68 5.50 5.59 55.62 98.29 71.04 50.01 12.16 0.96 1.78 52.51 98.13 68.41 50.00
D3R 6.99 80.99 12.86 56.47 99.98 72.17 51.39 32.87 75.97 45.89 61.47 78.52 68.96 79.95
PatchAD (Ours) 89.48(2.08) 31.22(0.74) 46.28(1.09) 73.07(0.41) 86.07(0.28) 79.04(0.12) 79.89(0.55) 85.42(0.98) 70.34(0.18) 77.15(0.48) 73.62(0.52) 84.65(0.22) 78.75(0.20) 91.19(0.12)
Dataset PSM SWAN
Metric P R F1_ori Aff-Pre AFF-Rec Aff-F1 AUC P R F1_ori Aff-Pre AFF-Rec Aff-F1 AUC
Deep SVDD 28.59 99.91 44.46 53.92 99.98 70.06 78.09 41.74 85.03 55.99 56.84 89.93 69.66 75.32
USAD 32.36 92.17 47.90 55.30 98.71 70.88 64.53 39.73 77.94 52.63 55.26 68.62 61.22 66.76
COUTA 100.00 0.04 0.08 100.00 5.12 9.75 72.45 62.96 0.09 0.17 69.25 0.10 0.19 71.38
TranAD 35.90 54.23 43.20 60.88 92.09 73.30 65.22 38.84 78.29 51.93 57.86 90.18 70.49 70.00
TimesNet 64.85 1.27 2.49 77.84 67.23 72.15 58.74 38.63 48.89 43.16 60.75 81.29 69.53 53.62
AnomTrans 26.05 84.32 39.81 55.35 80.28 50.09 52.18 31.19 79.87 44.86 58.45 9.49 0.54 45.38
DCdetector 27.05 19.59 22.72 52.33 76.95 65.88 49.62 32.45 98.62 48.83 38.73 3.13 7.50 50.35
GPT2-Adapter 28.31 46.64 35.24 54.33 95.16 69.17 50.74 40.00 52.05 45.23 61.17 69.55 65.09 58.77
D3R 28.96 93.65 44.24 53.84 100.00 69.99 60.64 36.12 78.38 49.45 57.70 96.06 72.10 62.54
PatchAD (Ours) 47.13(0.44) 49.32(0.97) 48.20(0.64) 63.52(0.82) 96.84(1.21) 76.72(0.19) 64.74(0.01) 51.76(0.18) 65.31(0.47) 57.75(0.08) 66.48(0.80) 79.39(1.54) 72.36(0.20) 72.91(0.02)
In Table 7, Aff-P and Aff-R are the precision and recall of affiliation metric [51], respectively. R_A_R and R_A_P are Range-AUC-ROC and Range-AUC-PR [52], respectively. V_ROC and V_RR are volumes under the surfaces created based on ROC curve and PR curve [52], respectively. Besides, we give more results in table 8, and the results from more real-world multivariate time-series datasets show that our proposed PatchAD has good robustness and performance.
Table 7: Multi-metrics results on all real-world multivariate datasets. All results are in %, the best in Bold, and the second in underlined.
@@
Table 8: More results on all real-world multivariate datasets. All results are in %, the best in Bold, and the second in underlined.
@@
Appendix H Limitations
Although PatchAD achieves optimal results on multiple datasets, some limitations still need to be addressed. In order to ensure the optimization of different metrics, subtle adjustments to potential parameters are necessary. For instance, selecting an appropriate anomaly threshold is crucial to balance precision and recall. Developing an automated threshold adjustment scheme is a potential avenue for future research. Time series anomaly detection employs a wide range of evaluation metrics, each with its own emphasis, leading to potential biases in evaluating different algorithms. In the future, we aim to propose a unified set of evaluation metrics that consider multiple factors, contributing to the advancement of the community.
Appendix I Broader Impacts
Time series anomaly detection has significant academic implications and broader impacts in various domains. In finance, it aids in detecting fraudulent activities like credit card fraud and insider trading, safeguarding financial systems and preventing losses. In healthcare, it aids in early disease detection and monitoring, enabling timely interventions and improved patient outcomes. Overall, time series anomaly detection improves decision-making, risk mitigation, and system performance across multiple domains.